OpenStudy (anonymous):

Solve the quadratic equation by factoring 12x^2-3x-9=0

OpenStudy (anonymous):

12x^2 - 12x + 9x - 9 = 0 12x (x - 1) + 9 (x - 1) =0 (12x + 9) (x - 1) = 0 12x + 9 = 0 x - 1 = 0 x = -9/12 x = 1

OpenStudy (anonymous):

Thank You.

OpenStudy (anonymous):

another way : factor the exact number for 12,3,9 which gives u 3 3(4x^2-x-3)=0 3(x-1)(4x+3)=0 3=0 x-1=0 >>> x=1 4x+3=0 >>> 4x=-3 >>> x=-3/4